a computer store sells two models of monitors. model a sells for $355 and model b sells for $465. if a total of 120 monitors were sold for $47880, how many of each type were sold?

Answer by ikleyn(52781)

.
a computer store sells two models of monitors. model a sells for $355
and model b sells for $465. if a total of 120 monitors were sold for $47880,
how many of each type were sold?
~~~~~~~~~~~~~~~~~

Let x be the number of the $465 panels;

then 120-x is the number of the $355 panels.


Write the total revenue equation

    465x + 355*(120-x) = 47880.


Simplify and find x

    465x + 355*120 - 355x = 47880

         110x             = 47880 - 355*120

         110x             =      5280

            x             =      5280/110 = 48.


ANSWER.  48 monitors at $465 and the rest, 120 - 48 = 72 monitors at $355.

Solved.